Thursday, 20 August

Plenary Lectures

9:30 – 10:30  Auditorium Maximum H 105

Dusa McDuff, SUNY, Stony Brook, USA Fibrations in Symplectic Topology (I: p.16)

Chair: Yakov Eliashberg, Stanford University, Stanford, USA

11:00 – 12:00  Auditorium Maximum H 105

I.G. Macdonald, Queen Mary and Westfield College, University of London, London, England Constant Term Identities, Orthogonal Polynomials, and Affine Hecke Algebras (I: p.12)

Chair: Richard Borcherds, University of Cambridge, Cambridge, UK

ICM’98 – Program  1